Find the area of the circle with the given circumference.
Use 3.14 for pi.
C=31.4 ft

Answer by stanbon(75887)

Find the area of the circle with the given circumference.
Use 3.14 for pi.
C=31.4 ft
-------------
C = 2(pi)r
31.4 = 2(3.14)r
r = 5
---
Area = (3.14)(5^2) = 78.5 sq. ft.
==================================
Cheers,
Stan H.